# A Language Anchor-Guided Method for Robust Noisy Domain Generalization

### v1

@@ 1 Introduction
Domain Generalization (DG) has emerged as a pivotal algorithm in machine learning, aiming to develop models that can maintain high performance on previously unseen environments—or domains. Traditional methods often assume that training and test data share the same distribution, yet in real-world scenarios, there is frequently a substantial shift between these distributions [1]. This phenomenon, widely referred to as domain shift, can cause severe performance degradation in tasks spanning computer vision, natural language processing, and medical image analysis [2]. As shown in Figure 1(a)(b), even within the same class label, the distribution of feature representations can vary considerably. This variation may stem from differences in image conditions, such as lighting, pose, or background, as well as other domain-specific factors. Such intra-class variability poses a significant challenge for models, which must learn to extract invariant features that capture the true semantic essence of the class while ignoring irrelevant variations. In response, DG task has been proposed to enable robust model behavior without further fine-tuning, thereby facilitating broader applicability in diverse domains [3, 4]. Over time, researchers have explored various ways to align or unify source and target domains, including strategies that learn domain-invariant features through multi-task autoencoders, augmented architecture, or adversarial alignment  [5, 6, 7].
@@
One persistent obstacle in DG is the prevalence of spurious correlations in deep neural networks, where models inadvertently link target labels to irrelevant, domain-specific features [8, 9, 10]. For instance, a classifier tasked with recognizing animals might rely on particular backgrounds for prediction, causing it to misclassify animals photographed in unfamiliar settings [11]. Such spurious correlations frequently arise because modern networks have the capacity to overfit even low-level noise, thereby anchoring their decisions on non-causal cues [12]. To address these pitfalls, researchers have proposed a variety of techniques, from causal inference algorithms like Invariant Risk Minimization (IRM) [8, 13] to methods specifically designed to disentangle spurious and invariant features [14]. For example, Wu et al. [15] introduced the Discover and Cure (DISC) strategy, which detects suspicious features and systematically mitigates their impact by leveraging interpretable domain knowledge. The development of methods rooted in causal representation learning also arises to mitigate the reliance on spurious correlations [16, 17, 18]. Despite these advances, achieving robust generalization under significant domain shift and considerable label noise persists as an open research challenge that necessitates more comprehensive methods and better theoretical insights.
@@
Existing domain generalization (DG) approaches have predominantly focused on aligning feature distributions or disentangling spurious factors using data-driven methods [2]. However, these techniques often struggle when label noise is prevalent or when the data lack clear domain-invariant cues [19], leading to rapid degradation in network performance as noise intensity increases. This challenge is further exacerbated in large, overparameterized networks that are prone to memorizing spurious correlations, thereby amplifying errors. Indeed, estimates suggest that 8% to 38.5% of real-world data may suffer from label corruption, undermining the generalizability of models that fail to distinguish genuine signals from incidental patterns [20]. Although some DG algorithms inherently exhibit robustness to label noise, in general cases they lack performance superiority compared to empirical risk minimization (ERM) baseline [10]. Consequently, understanding how to boost robustness against both domain shifts and noisy labels remains a pressing goal in building truly reliable artificial intelligence systems.
@@
In light of these challenges, we consider an alternative approach that focuses on introducing novel sources of features from diverse perspectives to guide the learning pipeline. An emerging line of research has investigated the role of external knowledge, which can guide models toward more semantically grounded representations [21, 22]. Some studies incorporate domain knowledge directly into the learning pipeline, for instance by embedding specialized scientific insights into feature extraction or loss design [23]. In other cases, researchers have shown that domain knowledge can significantly enhance model interpretability and robustness—for example, in laser-induced breakdown spectroscopy, where integrated expert knowledge improved quantification performance [24]. Another promising direction involves multi-modal alignment, in which semantic text descriptions serve as stable anchors that help networks transcend purely visual or sensor-driven biases [25]. Models like CLIP [26] exemplify this principle by projecting both text and images into a shared embedding space, improving generalization across dissimilar environments. Yet critical barriers remain regarding how to ensure that the guidance remains valid across diverse domains and how to structure knowledge encoding effectively to align with modern deep learning architectures [22].
@@
  • 2.
    We introduce the concept of NLP anchors derived from large-scale language models (e.g., CLIP) to provide domain-invariant, semantically rich feature constraints, significantly improving model robustness. Furthermore, we propose a novel weighted loss function that dynamically adjusts the contribution of each sample based on its distance to the corresponding NLP anchor. By assigning higher importance to samples that are closer to the learned semantic representations and lower importance to outliers, our approach effectively enhances feature alignment, leading to improved generalization across unseen domains.
  • 3.
@@
The remainder of the paper is organized as follows: In Section 2, we review the related work. Section 3 details our proposed method and addresses each of the key steps. In Section 4, we present experimental details along with an extensive analysis of the implementation and parameter choices. Section 5 discusses the limitations of our approach and potential strategies for addressing them. Finally, Section 6 concludes the paper and outlines directions for future research.
@@ 2.1 Domain Generalization
Domain generalization (DG) deals with the challenge of training models that can perform well on previously unseen domains, without access to labeled examples from those domains at training time. A significant challenge in DG is the presence of spurious correlations in deep neural networks, which hinder generalization across diverse settings [8, 9, 27, 28]. These correlations cause models to rely on domain-specific artifacts rather than truly invariant features, leading to performance degradation when deployed in unseen domains. Early methods focused on learning domain-invariant representations through shallow or deep feature alignments, aiming to eliminate domain-specific information while preserving task-relevant features. For instance, Muandet et al. [3] propose Domain-Invariant Component Analysis (DICA), a kernel-based method to learn invariant features by minimizing the discrepancy across source domains. Ghifary et al. [5] use multi-task autoencoders to learn generic representations.
@@
Beyond augmentation, several methods focus on regularization to improve generalization. Balaji et al. [33] introduced a regularization algorithm by learning a regularizer modeling the objective of DG that a model trained on one domain to generalize effectively. Dou et al. [34] utilized episodic training to simulate domain shifts and improve generalization. Zhou et al. [35] introduced a regularization approach based on domain-adaptive ensemble learning to enhance a model’s ability to generalize across different domains. Furthermore, recent research has revisited the issue of spurious correlations in DG, proposing methods to mitigate their impact. Li et al. [36] explored building a structural causal model for representation learning to address spurious correlations. Ma et al. [37] introduced FedCD, a federated domain generalization algorithm that employs a spurious correlation intervener for self-supervised feature intervention and a risk extrapolation aggregation strategy to reduce reliance on misleading shortcuts and boost performance on unseen domains. These studies underscore the importance of addressing spurious correlations to enhance DG.
@@ 2.2 Learning under noisy labels
2.3 NLP-Based Embeddings and External Knowledge
Concurrent to the advances in DG, there has been a surge of interest in leveraging NLP-based embeddings within computer vision and multimodal tasks. Early work such as DeViSE [46] demonstrated the potential of semantic embedding spaces derived from word vectors for zero-shot classification, effectively bridging visual and textual representations. More recently, contrastive vision-language models like CLIP [26] have illustrated that aligning images with corresponding textual descriptions in a shared embedding space can yield robust and generalizable features for zero-shot recognition. Beyond zero-shot applications, the integration of NLP embeddings in broader tasks such as image captioning, visual question answering, and cross-modal retrieval has significantly improved performance, suggesting that textual embeddings provide crucial semantic cues often absent in purely vision-based models [47, 48]. The integration of external knowledge has been shown to enhance the performance of large language models on domain-specific tasks. Zhang et al. [49] evaluated the fusion of external and parametric knowledge in large language models, highlighting the benefits of such integration.However, most of these approaches focus on either downstream supervised tasks or zero-shot learning scenarios, where textual labels serve as direct supervision signals. In contrast, domain generalization settings require learning representations that are intrinsically resilient to distribution shifts, without having direct access to data from target domains.
Our proposed algorithm, $A^{3}W$, leverages NLP anchors—semantic representations derived from large-scale language models—in a DG context to guide feature extraction and alignment. By intertwining visual signals with text-derived semantics, $A^{3}W$ effectively imposes additional structure on the learned representations, mitigating domain-specific idiosyncrasies.
@@ 3.1 Problem Formulation
Noisy Domain Generalization:
The proposed method aims to generalize to unseen domains while preventing overfitting to noisy data. In this setting, the training dataset $\mathcal{S}_{tr}$ is contaminated with label noise. We assume that the testing dataset $\mathcal{S}_{te}$ is drawn from a clean distribution with labels $\mathcal{Y}_{te}$. Our objective is to train a model on the noisy training dataset $\mathcal{S}_{tr}$ that learns robust and domain-invariant representations, thereby minimizing the classification error on the clean target domain $\mathcal{D}_{te}$.
@@ 3.3 Algorithm for $A^{3}W$
2) Main Model Update: we update the featurizer and classifier to improve classification performance using the combined loss of the weighted alignment and cross-entropy. During inference, the exponential moving average (EMA) network, initialized as a deep copy of the primary network, is employed to produce predictions, as it generally yields more stable and robust outputs compared to the primary network.
@@ 3.6 Weighted Loss
Alignment Loss
One way to enforce that the features align with the corresponding NLP anchor is to minimize the negative cosine similarity loss between the mapped feature representation $Proj_{y_{i}}(f(\mathbf{x}_{i}))$ and the fixed NLP anchor $\mathbf{a}_{y_{i}}$. This loss is defined as
@@
Minimizing $\mathcal{L}_{\text{anchor}}$ effectively maximizes the cosine similarity, ensuring that the learned features are directionally aligned with the semantic anchors and thus promoting semantically meaningful and domain-invariant representations.
Weight Computation
To prioritize high-confidence training instances, we assign importance weights to samples based on their similarity to the NLP anchors. We introduce a temperature parameter $\tau$ to adjust the sharpness of the softmax weighting distribution—higher values of $\tau$ yield a more peaked distribution that emphasizes strongly aligned samples, whereas lower values produce a softer weighting. Specifically, we first compute the softmax weights over the alignment costs:
@@
Overall Loss
To maintain classification accuracy while leveraging the selective sampling, we incorporate a weighted cross-entropy loss, where samples with higher $w_{i}$ values contribute more significantly to the overall loss. The final loss function combines the alignment loss and the weighted cross-entropy loss:
@@
$\mathcal{L}\;=\;\sum_{i=1}^{N}w_{i}\Biggl{[}\lambda\,\Bigl{(}-\cos\bigl{(}Proj_{y_{i}}(f(\mathbf{x}_{i})),\,\mathbf{a}_{y_{i}}\bigr{)}\Bigr{)}\;+\;\ell_{CE}\Bigl{(}g\bigl{(}f(\mathbf{x}_{i})\bigr{)},\,y_{i}\Bigr{)}\Biggr{]}$ (13)
From this formulation, the weight term $w_{i}$ acts as an importance factor that regulates the influence of each sample on the overall loss. This regularization mechanism ensures that both the alignment loss and the classification loss emphasize reliable, high-confidence data. By down-weighting the contributions of samples that are likely to be noisy or out-of-distribution, the weight term effectively mitigates the impact of such noise on the model’s training, leading to more robust feature learning and improved classification accuracy.
3.6.1 Iterative Update
@@ 3.7 Theoretical Insights
A uniformly bounded gradient implies that local changes in $f(x)$ away from $\mathbf{a}_{c}$ incur non-negligible cost. Hence, any hypothesis that aligns poorly with anchors sees a high penalty, effectively restricting the feasible region of $\mathcal{H}$. ∎
Lemma 3.1 shows that anchor alignment behaves like a regularizer, steering the network away from memorizing noise-laden features.
To further enhance robustness, we employ a continuous weighting scheme (Eq. (11)) that assigns higher importance to samples with strong semantic alignment. This reweighting mechanism effectively adjusts the empirical risk, as captured in our overall loss function (Eq. (12)), so that samples likely to be correctly labeled have a greater influence during training.
@@
Theorem 3.1 indicates that our weighting strategy can mitigate the detrimental effects of noise, enabling the model to approximate the true (clean) distribution more closely.
These results not only demonstrate the efficacy of our reweighting strategy in handling noisy labels but also motivate a further examination of how our approach reduces the discrepancy between noisy and clean distributions. Let $P$ and $Q$ be the distributions of the source and target domains, respectively, with label noise in $P$. Many domain generalization results [53] rely on bounding a distributional divergence $\mathrm{div}(P,Q)$. Noise can inflate this divergence by altering label proportions or feature-label mappings. However, semantic alignment and selective reweighting encourage the model to focus on consistent semantic cues, thereby lowering the effective divergence to the clean distribution.
@@
4 Experiments
4.1 Experimental Setup and Datasets Preprocessing
We conducted our experiments on a server with 10 NVIDIA Quadro RTX 6000 24G GPUs. We extensively evaluate our method on domain generalization with noisy data using four benchmark datasets. For more details, please refer to their original publications. PACS [54] is a 7-class image classification dataset that spans four distinct domains (Photo, Art Painting, Cartoon, and Sketch). Renowned for its diverse artistic styles, PACS offers a challenging testbed for robust representation learning. VLCS [47] comprises 5 classes drawn from four domains (Caltech101, LabelMe, SUN09, and VOC2007). Each domain originates from a different source, resulting in significant variability in image characteristics and presenting a broad generalization challenge with both natural and scene-centric images. Office-Home [55] is a 65-class dataset designed to capture common objects in everyday office and home environments. Its wide range of object categories and the substantial variation in style and background make it a rigorous benchmark for domain generalization. SVIRO [56] is a synthetic dataset focused on vehicle interiors. It contains 25,000 images from 10 distinct vehicle interior environments and features 7 occupant classes. SVIRO provides a challenging scenario for domain generalization, especially in handling variations in interior design and occupancy. DomainNet  [57] is a large-scale dataset spanning six domains (Clipart, Infograph, Painting, Quickdraw, Real, Sketch) with 345 classes. It encompasses a broad range of styles and object categories, capturing significant distribution shifts across these domains. Due to computation resource limitation, we only used the first four domains in SVIRO and 15 classes per domain for DomainNet. More information is provided in  1.
Table 1: Dataset Information for Domain Generalization Benchmarks

### v2

@@ 1 Introduction
Domain Generalization (DG) has emerged as a pivotal algorithm in machine learning, aiming to develop models that can maintain high performance on previously unseen environments—or domains. Traditional methods often assume that training and test data share the same distribution, yet in real-world scenarios, there is frequently a substantial shift between these distributions. This phenomenon, widely referred to as domain shift, can cause severe performance degradation in tasks spanning computer vision, natural language processing, and medical image analysis [1]. As shown in Figure 1(a)(b), even within the same class label, the distribution of feature representations can vary considerably. This variation may stem from differences in image acquisition conditions—such as lighting variations, changes in pose, or complex background environments—and even from more subtle domain-specific factors like sensor noise or camera calibration differences. Such intra-class variability poses a significant challenge for developing accurate and adaptable models, which must learn to extract invariant features that capture the true semantic essence of the class while ignoring irrelevant variations. In response, DG task has been proposed to enable robust model behavior without further fine-tuning, thereby facilitating broader applicability in diverse domains [2, 3]. Over time, researchers have explored various ways to align or unify source and target domains, including strategies that learn domain-invariant features through multi-task autoencoders, augmented architecture, or adversarial alignment [4, 5, 6].
@@
One persistent obstacle in DG is the prevalence of spurious correlations in deep neural networks, where models inadvertently link target labels to irrelevant, domain-specific features [7, 8]. For instance, a classifier tasked with recognizing animals might rely on particular backgrounds for prediction, causing it to misclassify animals photographed in unfamiliar settings [9]. Such spurious correlations frequently arise because modern networks have the capacity to overfit even low-level noise, thereby anchoring their decisions on non-causal cues [10]. To address these pitfalls, researchers have proposed a variety of techniques, from causal inference algorithms like Invariant Risk Minimization (IRM) [7] to methods specifically designed to disentangle spurious and invariant features [11]. For example, Wu et al. [12] introduced the Discover and Cure (DISC) strategy, which detects suspicious features and systematically mitigates their impact by leveraging interpretable domain knowledge. The development of methods rooted in causal representation learning also arises to mitigate the reliance on spurious correlations [13]. Despite these advances, achieving robust generalization under significant domain shift and considerable label noise persists as an open research challenge that necessitates more comprehensive methods and better theoretical insights.
@@
Existing domain generalization (DG) approaches have predominantly focused on aligning feature distributions or disentangling spurious factors using data-driven methods [1]. While these techniques have paved the way for significant advances, they often encounter considerable difficulties when label noise is prevalent or when data lack distinct domain-invariant cues [14], leading to rapid degradation in network performance as noise intensity increases. This challenge is further exacerbated in large, overparameterized networks that are prone to memorizing spurious correlations, thereby amplifying errors. Indeed, estimates suggest that 8% to 38.5% of real-world data may suffer from label corruption, undermining the generalizability of models that fail to distinguish genuine signals from incidental patterns [15]. Although some DG algorithms inherently exhibit robustness to label noise, in general cases they lack performance superiority compared to empirical risk minimization (ERM) baseline [8]. Consequently, understanding how to boost robustness against both domain shifts and noisy labels remains a pressing goal in building truly reliable artificial intelligence systems.
@@
In light of these challenges, we consider an alternative approach that focuses on introducing novel sources of features from diverse perspectives to guide the learning pipeline. An emerging line of research has investigated the role of external knowledge, which can guide models toward more semantically grounded representations [16]. Some studies incorporate domain knowledge directly into the learning pipeline, for instance by embedding specialized scientific insights into feature extraction or loss design [17]. In other cases, researchers have shown that domain knowledge can significantly enhance model interpretability and robustness—for example, in laser-induced breakdown spectroscopy, where integrated expert knowledge improved quantification performance [18]. Another promising direction involves multi-modal alignment, in which semantic text descriptions serve as stable anchors that help networks transcend purely visual or sensor-driven biases [19]. Models like CLIP [20] exemplify this principle by projecting both text and images into a shared embedding space, improving generalization across dissimilar environments. Yet critical barriers remain regarding how to ensure that the guidance remains valid across diverse domains and how to structure knowledge encoding effectively to align with modern deep learning architectures [16].
@@
  • We introduce the concept of NLP anchors derived from large-scale language models (e.g., CLIP), which provide domain-invariant and semantically rich feature constraints that significantly improve model robustness. Furthermore, we propose a novel weighted loss function that dynamically adjusts each sample’s contribution based on its distance to the corresponding NLP anchor. By assigning higher importance to samples that are closer to the learned semantic representations and lower importance to outliers, our approach effectively enhances feature alignment, ultimately leading to improved generalization across unseen domains.
@@ 2.1 Domain Generalization
Domain generalization (DG) addresses the challenge of training models that can perform well on previously unseen domains, without access to labeled examples from those domains at training time. A significant challenge in DG is the presence of spurious correlations in deep neural networks, which hinder generalization across diverse settings [7]. These correlations arise when models inadvertently rely on domain-specific artifacts—features that are incidentally correlated with the target labels in the training data—rather than on the truly invariant properties that are essential for robust performance. Consequently, when models encounter data from an unseen domain, their overreliance on these spurious cues leads to performance degradation when deployed in unseen domains. Early methods focused on learning domain-invariant representations through shallow or deep feature alignments, aiming to eliminate domain-specific information while preserving task-relevant features. For instance, Muandet et al. [2] propose Domain-Invariant Component Analysis (DICA), a kernel-based method to learn invariant features by minimizing the discrepancy across source domains. Ghifary et al. [4] use multi-task autoencoders to learn generic representations.
@@
Beyond augmentation, several methods have turned their focus to regularization to improve generalization. Balaji et al. [25] introduced a regularization algorithm by learning a regularizer modeling the objective of DG that a model trained on one domain to generalize effectively. Dou et al. [26] utilized episodic training to simulate domain shifts and improve generalization. Zhou et al. [27] introduced a regularization approach based on domain-adaptive ensemble learning to enhance a model’s ability to generalize across different domains. Furthermore, recent research has revisited the issue of spurious correlations in DG, proposing methods to mitigate their impact. Li et al. [28] explored building a structural causal model for representation learning to address spurious correlations. Ma et al. [29] introduced FedCD, a federated domain generalization algorithm that employs a spurious correlation intervener for self-supervised feature intervention and a risk extrapolation aggregation strategy to reduce reliance on misleading shortcuts and boost performance on unseen domains. These studies underscore the importance of addressing spurious correlations to enhance DG.
@@ 3.1 Problem Formulation
Noisy Domain Generalization: The proposed method aims to generalize to unseen domains while preventing overfitting to noisy data. In this setting, the training dataset $\mathcal{S}_{tr}$ is contaminated with label noise. We assume that the testing dataset $\mathcal{S}_{te}$ is drawn from a clean distribution with labels $\mathcal{Y}_{te}$. Our objective is to train a model on the noisy training dataset $\mathcal{S}_{tr}$ that learns robust and domain-invariant representations, thereby minimizing the classification error on the clean target domain $\mathcal{D}_{te}$.
@@ 3.3 Algorithm for $A^{3}W$
2) Main Model Update: we subsequently update the featurizer and classifier to improve classification performance using the combined loss of the weighted alignment and cross-entropy. During inference, the exponential moving average (EMA) network, initialized as a deep copy of the primary network, is employed to produce predictions, as it generally yields more stable and robust outputs compared to the primary network.
@@ 3.6 Weighted Loss
Alignment Loss: One way to enforce that the features align with the corresponding NLP anchor is to minimize the negative cosine similarity loss between the mapped feature representation $Proj_{y_{i}}(f(\mathbf{x}_{i}))$ and the fixed NLP anchor $\mathbf{a}_{y_{i}}$. This loss is defined as
@@
By minimizing $\mathcal{L}_{\text{anchor}}$, we effectively maximizes the cosine similarity between the projected features and the NLP anchor. This process ensures that the learned features are directionally aligned with the semantic anchors, ultimately promoting representations that are both semantically meaningful and domain-invariant.
Weight Computation: To prioritize high-confidence training instances, we assign importance weights to samples based on their similarity to the NLP anchors. We introduce a temperature parameter $\tau$ to adjust the sharpness of the softmax weighting distribution—higher values of $\tau$ yield a more peaked distribution that emphasizes strongly aligned samples, whereas lower values produce a softer weighting. Specifically, we first compute the softmax weights over the alignment costs:
@@
Overall Loss: To maintain classification accuracy while leveraging the selective sampling, we incorporate a weighted cross-entropy loss, where samples with higher $w_{i}$ values contribute more significantly to the overall loss. The final loss function combines the alignment loss and the weighted cross-entropy loss:
@@
$\mathcal{L}\;=\;\sum_{i=1}^{N}w_{i}\Biggl{[}\lambda\,\Bigl{(}-\cos\bigl{(}Proj_{y_{i}}(f(\mathbf{x}_{i})),\,\mathbf{a}_{y_{i}}\bigr{)}\Bigr{)}\;+\;\ell_{CE}\Bigl{(}g\bigl{(}f(\mathbf{x}_{i})\bigr{)},\,y_{i}\Bigr{)}\Biggr{]}$
From this formulation, the weight term $w_{i}$ acts as an importance factor that regulates the influence of each sample on the overall loss. This regularization mechanism ensures that both the alignment loss and the classification loss emphasize reliable, high-confidence data. By down-weighting the contribution of samples that are likely noisy or out-of-distribution, the weight term effectively minimizes the adverse impact of noise during training. Consequently, this leads to more robust feature learning and improved classification accuracy.
3.7 Iterative Update
@@ 3.8 Theoretical Insights
A uniformly bounded gradient implies that local changes in $f(x)$ away from $\mathbf{a}_{c}$ incur non-negligible costs. Hence, any hypothesis that aligns poorly with anchors sees a high penalty, effectively restricting the feasible region of $\mathcal{H}$. ∎
Lemma 3.1 shows that anchor alignment behaves like a regularizer, steering the network away from memorizing noise-laden features. To further enhance robustness, we employ a continuous weighting scheme (Eq. (11)) that assigns higher importance to samples with strong semantic alignment. This reweighting mechanism effectively adjusts the empirical risk, as captured in our overall loss function (Eq. (12)), so that samples likely to be correctly labeled have a greater influence during training.
@@
Theorem 3.1 indicates that our weighting strategy can mitigate the detrimental effects of noise, enabling the model to approximate the true (clean) distribution more closely. These results not only demonstrate the efficacy of our reweighting strategy in handling noisy labels but also motivate a further examination of how our approach reduces the discrepancy between noisy and clean distributions. Let $P$ and $Q$ be the distributions of the source and target domains, respectively, with label noise in $P$. Many domain generalization results rely on bounding a distributional divergence $\mathrm{div}(P,Q)$. Noise can inflate this divergence by altering label proportions or feature-label mappings. However, semantic alignment and selective reweighting encourage the model to focus on consistent semantic cues, thereby lowering the effective divergence to the clean distribution.
@@
TABLE I: Dataset Information for Domain Generalization Benchmarks
@@
4 Experiments
4.1 Experimental Setup and Datasets Preprocessing
We conducted our experiments on a server with 10 NVIDIA Quadro RTX 6000 24G GPUs. We extensively evaluate our method on domain generalization with noisy data using four benchmark datasets. For more details, please refer to their original publications. PACS [32] is a 7-class image classification dataset that spans four distinct domains (Photo, Art Painting, Cartoon, and Sketch). Renowned for its diverse artistic styles, PACS offers a challenging testbed for robust representation learning. VLCS [30] comprises 5 classes drawn from four domains (Caltech101, LabelMe, SUN09, and VOC2007). Each domain originates from a different source, resulting in significant variability in image characteristics and presenting a broad generalization challenge with both natural and scene-centric images. Office-Home [33] is a 65-class dataset designed to capture common objects in everyday office and home environments. Its wide range of object categories and the substantial variation in style and background make it a rigorous benchmark for domain generalization. SVIRO [52] is a synthetic dataset focused on vehicle interiors. It contains 25,000 images from 10 distinct vehicle interior environments and features 7 occupant classes. SVIRO provides a challenging scenario for domain generalization, especially in handling variations in interior design and occupancy. DomainNet  [43] is a large-scale dataset spanning six domains (Clipart, Infograph, Painting, Quickdraw, Real, Sketch) with 345 classes. It encompasses a broad range of styles and object categories, capturing significant distribution shifts across these domains. Due to computation resource limitation, we only used the first four domains in SVIRO and 15 classes per domain for DomainNet. More information is provided in  I.